OpenStudy (anonymous):

find cos (alpha + Beta). sin alpha=5/13, alpha lies in Q1, and cos Beta=15/17 is also in Q1, I already have B= +/- 12, and A=+/-8, please help? :)

OpenStudy (dumbcow):

use sum identity for cosine \[\cos(a+b) = \cos(a) \cos(b) - \sin(a) \sin(b)\] so you need sin(b) and cos(a) for that use pythagorean identity \[\sin ^{2} \theta + \cos^{2} \theta = 1\] \[(\frac{5}{13})^{2} + \cos^{2} a = 1\] \[\cos a = \frac{12}{13}\] do same thing to find sin(b)

OpenStudy (anonymous):

I'm still pretty lost here :/

OpenStudy (dumbcow):

lets try triangle approach you know that sin(a) = 5/13 and cos(b) = 15/17 draw 2 triangles reflecting those trig ratios  solve for the missing sides of the triangles to get the other trig ratios does that make sense

OpenStudy (anonymous):

agh, these just dont make sense to me, im sorry

OpenStudy (dumbcow):

but you said you already got 12 and 8 ... obviously you used pythagorean thm to get these numbers right?

OpenStudy (dumbcow):